OpenStudy (anonymous):

Factor : x^2 - x - 3

OpenStudy (anonymous):

It needs to be in parenthesis. Like (x + ?)(x - ? )

OpenStudy (anonymous):

yes but first u need to know if it is factor-able in the first place!

OpenStudy (anonymous):

IT IS FRIGGIN' FACTORABLE.

OpenStudy (anonymous):

use the quadratic formula \[\left( -b \pm \sqrt{b ^{2}-4ac}\right)/2a\]

OpenStudy (anonymous):

I am a freshman in High SChool. I have no idea what you just said.

OpenStudy (apoorvk):

quadratic formula compare your quadratic expression to the standard equation \[ax^2 + bx +c\] then the values of x will be : \[x = [-b \pm \sqrt{b^2 - 4ac}]/2a\] now you can write them in their factor form

OpenStudy (anonymous):

hay girl quadratic formula is use to solve the quadratic equations like the one you have asked: x^2 - x - 3 now in this eq b=-1 a=1 c=-3 put these values in the quation you'll get you answer ok ?

OpenStudy (diyadiya):

@Lovelesstime what did you get?